[PRad] RC weekly report

gasparan at jlab.org gasparan at jlab.org
Sun May 8 23:37:39 EDT 2016

Ran Coordinator Report for the past week.

1)	Tagger electronics is updated. Communication to HV crates established.
All T-counters (except one) are powered up and looking good. All
E-counters (except few) are powered up and looking good (thanks to Eugene
and Nathan). Some work still remains on Epics part, planned for Monday.
2)	Good progress on the target part. All pumps are currently installed and
operational. Planned vacuum level is established in the target chamber.
Camber is connected to the downstream Vacuum tank.  Pumping down started
Sunday ~9 pm. Hydrogen tests on Monday.
3)	Vacuum tank is installed in the beam line. Pumps are installed and
~10^-5 level vacuum was established in the Tank. It is ready for the
4)	Trigger scintillator detectors are installed on the exit window of the
Vacuum tank. All cabling work is done. Cosmic tests are performed. Ready
for the experiment.
5)	GEM detectors are installed on HyCal. The shelf for the GEM electronics
is installed underneath of HyCal. Cabling work is done. Gas supply
connections are established. Ar+CO2 gas circulation for both chambers will
start on Monday.
6)	 HyCal is fully repaired on Saturday, ready for the experiment. Cooling
down will start on Monday.
7)	Good progress on DAQ system. Still some remaining issues with the DAQ
speed (~3 KHz level), there are some ideas. Work will be continued on
8)	Tests on the on-line programs started on Saturday/Sunday.  This work
will continue from Monday.

Overall, a very productive work is performed in the past week.
Estimated experiment readiness level:   ~80%
Many thanks to all students, postdocs, Target group and Hall B members.


More information about the PRad mailing list